# Zero-Shot Generalization during Instruction Tuning: Insights from Similarity and Granularity

### v1

@@ Abstract
Understanding alignment techniques begins with comprehending zero-shot generalization brought by instruction tuning, but little of the mechanism has been understood. Existing work has largely been confined to the task level, without considering that tasks are artificially defined and, to LLMs, merely consist of tokens and representations. This line of research has been limited to examining transfer between tasks from a task-pair perspective, with few studies focusing on understanding zero-shot generalization from the perspective of the data itself. To bridge this gap, we first demonstrate through multiple metrics that zero-shot generalization during instruction tuning happens very early. Next, we investigate the facilitation of zero-shot generalization from both data similarity and granularity perspectives, confirming that encountering highly similar and fine-grained training data earlier during instruction tuning, without the constraints of defined “tasks”, enables better generalization. Finally, we propose a more grounded training data arrangement method, Test-centric Multi-turn Arrangement, and show its effectiveness in promoting continual learning and further loss reduction. For the first time, we show that zero-shot generalization during instruction tuning is a form of similarity-based generalization between training and test data at the instance level. We hope our analysis will advance the understanding of zero-shot generalization during instruction tuning and contribute to the development of more aligned LLMs. [^1] Our code is released at https://github.com/HBX-hbx/dynamics_of_zero-shot_generalization
@@ 1 Introduction
The extraordinariness of large language models (LLMs) was originally brought by the zero-shot generalization of instruction tuning [2]. At the early stage, studies have found that when diverse prompts are added to the inputs of traditional natural language processing (NLP) tasks and fed into the model for instruction tuning, the model can generalize to tasks it has never encountered before [5, 24, 37, 47, 48]. To date, instruction tuning [5, 37, 48] has become a crucial phase in training large language models, often preceding methods that incorporate preference data. In the meantime, the concept of “task” is also becoming increasingly blurred. Researchers are no longer constructing instruction data in the ways traditional NLP tasks dictate, but rather, they hope these tasks will be as close to reality and as diverse as possible [4, 11, 35, 41, 54].
@@
Although nearly all LLMs benefit from the zero-shot generalization brought about by instruction tuning, the in-depth and fine-grained research on this phenomenon is still insufficient. Particularly, there are few accurate and comprehensive conclusions about when and in what form it occurs, and how the data could influence it. A line of research focuses on understanding the relationships in task-pair transfer [40, 45, 52], suggesting that not all tasks contribute positively to zero-shot generalization; some tasks may even result in negative transfer effects [21, 26, 56]. The main limitation of the existing works is that they are confined within the concept of “task”. Whether they train on one task and then evaluate on another [21, 56], or simply calculate intermediate task [45] or instruction [22] transfer scores, all these efforts to select the most informative tasks to promote zero-shot generalization are restricted to the task-based framework. This approach is based on the premise that human-defined “tasks” and even “categories” are sufficiently reasonable. However, this is often not the case as gaps exist regarding how humans and LLMs perceive the instruction tuning data. To this end, we strive to break free from the task-level framework and explore “generalization” at a more fundamental level.
@@
In this paper, we conduct a comprehensive investigation of the zero-shot generalization during instruction tuning and attempt to answer some critical questions: In instruction tuning, i) when does zero-shot generalization occur? ii) How can we more accurately understand the role of data in zero-shot generalization? iii) How can we effectively promote zero-shot generalization?
To answer the first research question, we attempt to pinpoint the timing of zero-shot generalization during instruction tuning in. In line with existing works [3, 16, 53, 55], we discover that zero-shot generalization occurs extremely early regardless of the metrics we apply for measurement. This indicates that LLM’s instruction-following ability could be unlocked by leveraging merely a few sample training data. In addition, we find loss as a stable and fair indicator of zero-shot generalization during our measurement. (Section 2)
@@
To further investigate why zero-shot generalization occurs early, we explore the impact brought by different data permutations, demonstrating that training data itself plays an important role during the early stage of instruction tuning. Diving deeper, we identify two perspectives in understanding this role: similarity and granularity. From a similarity perspective, we discover that the model’s generalization is not truly “zero-shot”, as a high resemblance between the test and training data distributions could significantly impact generalization. To quantify the similarity, we define Weighted Similarity Distance (WSD), taking into account both the average and minimum cosine similarity (Cosine-Avg and Cosine-Min) between the training data seen by the model and the test data. Our experiments reveal that encountering data with high similarity (low WSD) early during instruction tuning enhances zero-shot generalization compared to encountering such data later. From a granularity perspective, we disclose that artificially defined “tasks” are no longer suitable for measuring generalization. Instead, the similarity between training data and test data serves as a more essential indicator. By treating all data points equally and considering them at the instance level without the constraints of pre-defined “tasks,” we can better promote zero-shot generalization. (Section 3)
Though the WSD measure shows effectiveness in promoting zero-shot generalization, it still suffers from inherent limitations in that the individual role of each test data point cannot be distinguished, as the whole test set is considered holistically during distance calculation. To this end, we propose the Test-centric Multi-turn Arrangement (TMA). TMA arranges training data according to each individual test data point and seeks to enhance zero-shot generalization regardless of the presence of predefined “tasks.” We demonstrate through experiments the effectiveness of TMA, and unveil that early access to data of higher similarity during instruction tuning can facilitate continual learning and further loss reduction. (Section 4)
@@
  • We show that zero-shot generalization occurs at the very early stage during instruction tuning, while loss serves as a reasonable and suitable metric to measure zero-shot generalization due to its stability and fairness across datasets.
  • We identify similarity and granularity as two perspectives to gain a deeper understanding of zero-shot generalization, revealing that encountering highly similar and fine-grained training data earlier during instruction tuning, enables better generalization.
  • We propose the Test-centric Multi-turn Arrangement, a more grounded training data arrangement method, and show that accessing high-similarity data during instruction tuning can facilitate continual learning and further loss reduction.
@@ 2 Positioning Zero-Shot Generalization
Early research shows that instruction tuning, which applies to various NLP tasks formatted with instructions, can generalize to various unseen tasks. However, most studies [5, 19, 24] focus on integrating diverse tasks or instruction templates, using human-generated or synthetic data, or exploring different fine-tuning strategies, while few studies address when zero-shot generalization actually occurs. Therefore, to bridge this gap, we seek to identify the positioning of zero-shot generalization. We begin by giving a formalization of zero-shot generalization.
Formalization. In multi-task scenarios, zero-shot generalization refers to the ability to perform effectively on unseen tasks, while only trained on a subset of tasks. For each task $T\in\mathcal{D}_{\text{Seen}}\cup\mathcal{D}_{\text{Unseen}}$, there exists an instructional description $I^{T}$, as well as several instances, where each instance is composed of an input $x^{T}_{i}$ and an output $y^{T}_{i}$. We define a model $M$ as capable of generalization on unseen tasks if, after training on every $T\in\mathcal{D}_{\text{Seen}}$, given an unseen task $T\in\mathcal{D}_{\text{Unseen}}$, and for any $(x^{T}_{i},y^{T}_{i})$, the model’s output $\hat{y}=M(I^{T},x^{T}_{i})$ and the label $y^{T}_{i}$ achieve a score surpassing a certain threshold, with regard to the selected metrics, indicating successful generalization on the task $T$.
2.1 Occurrence of Zero-Shot Generalization in Instruction Tuning
@@
Data and Settings. We utilize three multi-task datasets, namely Natural Instructions V2 (NIV2) [47], Public Pool of Prompts (P3) [37], and Flan-mini [14], for our analysis. For NIV2, we utilize the default track and default training-test split for instruction tuning and evaluation. For P3, we employ training and test tasks consistent with the vanilla T0 model [^2] https://huggingface.co/bigscience/T0pp. For Flan-mini, we randomly partition the training and test tasks. And we choose pre-trained LLaMA-2-7B [43] as our base model. For other details including dataset details, hyper-parameters, data concatenation, and prompt template, please refer to Appendix A. All subsequent experiments are based on this setup, where we adopt a fine-grained approach to save a series of full-parameter fine-tuning checkpoints, evaluating each on the test set to observe the results on the specified metric.
Metrics. Evaluation metrics are crucial for assessing generalization. For datasets including P3, Flan-mini, and OPT-IML [19], Exact-Match is commonly applied due to its simplicity. NIV2 additionally incorporates ROUGE-1 and ROUGE-L as metrics. Besides, in reinforcement learning scenarios, the reward model (RM) plays a vital role [7, 51], serving as a proxy for human preferences. This makes the RM score also a reasonable metric to reflect zero-shot generalization.
To comprehensively study the impact of metrics, we apply all the metrics listed above, including Exact-Match, ROUGE-1, ROUGE-L, and RM score, in our preliminary study. Especially, for the RM score, we use UltraRM-13B [7] as the reward model.
@@
Results. We demonstrate that zero-shot generalization occurs at a very early stage during instruction tuning. As depicted in the left plot of Figure 1, using ROUGE-1, ROUGE-L, and Exact-Match as metrics, the ROUGE scores rise from approximately 15 to over 35 in just about 10 training steps, indicating significant generalization with only around 160 training samples. In the middle plot, the RM score exhibits a similar trend, stabilizing around 50 steps across all three datasets.
It is also noteworthy that all these metrics exhibit certain problems: ROUGE-1, ROUGE-L, and Exact-Match as metrics entail the resulting curves being seriously unstable, while the RM score for NIV2 is significantly higher than those for the other two datasets, indicating a certain bias induced. We provide a more detailed discussion in Section A.5. Therefore, we seek to apply a more reasonable metric as an indicator to evaluate zero-shot generalization.
2.2 Loss as the Measurement for Zero-Shot Generalization
Loss is commonly applied across various model pre-training and fine-tuning scenarios. For example, the scaling law [6, 17, 20] entails predicting loss based on model parameter count and dataset size, unsupervised learning uses cross-entropy loss to quantify the difference between probability distribution, etc. Recent studies have also delved into understanding emergent abilities from the perspective of loss, indicating that when pre-training loss drops below a specific threshold, the model can perform well on downstream tasks [12]. All these measures suggest loss to be a promising metric for evaluating generalization on unseen tasks. However, no currently established practice applies it to measure zero-shot generalization. To bridge this gap, we comprehensively study and justify that:
{bclogo}
@@ 2.2 Loss as the Measurement for Zero-Shot Generalization
Data and Settings. We use the same dataset as in the previous experiment. We randomly sample 120 data points from all unseen tasks and generate outputs using a series of instruction-tuned checkpoints. We then calculate the average cross-entropy loss against the corresponding labels within each step. This also entails the same setting as in the previous experiment. Please refer to Section A.4 for more details.
Results. Zero-shot generalization similarly occurs at an early stage of instruction tuning with loss as the evaluation metric. As shown in the right plot of Figure 1, all three datasets reach their lowest points in terms of loss within less than 50 steps, which further strengthens our conclusion that zero-shot generalization occurs early.
@@
3 Facilitating Zero-Shot Generalization
Acknowledging the importance of metrics in measuring the positioning of zero-shot generalization, we next seek to investigate why generalization occurs at an early stage and what role training data plays during this phase. Our initial focus lies in the analysis of how training data permutations affect zero-shot generalization. Then, we investigate the facilitation of zero-shot generalization from both data similarity and granularity perspectives.
3.1 Effect of Training Data Permutations
@@
We have demonstrated that zero-shot generalization occurs at the early stage during instruction tuning. However, the reasons why models can achieve noteworthy generalization during this phase await investigation. The model receives only a limited amount of supervised data during this phase. Therefore, despite the scarcity, these data ought to play a significant role in facilitating generalization. With this intuition, we investigate the impact of exposure to different training data permutations during the early stage of instruction tuning.
Data and Settings. We apply 1600 Flan-mini training tasks to get a series of instruction-tuned checkpoints and evaluate them on various Flan-mini unseen tasks. Please refer to Section B.1 for more details. Specifically, we examine the following three training data permutations:
  • Round-robin: We employ round-robin scheduling to disperse training tasks as much as possible. Consequently, the model iterates over one data point from each task every 100 training steps.
  • Cluster: We arrange all data from each task together, resulting in task-level clusters throughout the entire training dataset. Consequently, the model trains on 8 new tasks every 10 steps.
@@
Results. Different training data permutations entail different loss curve patterns. As shown in Figure 2, the pattern of random and round-robin scheduling is similar due to round-robin being an extreme form of data shuffling. However, cluster scheduling differs largely from both. At certain steps during the instruction tuning, there exists a sudden decrease in average loss across different test tasks. This further highlights that leveraging a relatively small amount of data may induce a substantial drop in the average loss, thus causing a great impact on zero-shot generalization.
3.2 Zero-Shot Generalization through Data Similarity and Granularity
We have observed that training data permutation could lead to significant changes in the loss curve and that the timing of presence for certain data may greatly facilitate generalization on unseen tasks. With these findings, we naturally ask what is the best permutation that facilitates zero-shot generalization the most, and how to arrange these “certain data” that promote early generalization. In the following sections, we seek to address these questions through two perspectives: similarity and granularity.
3.2.1 Enhancing Zero-Shot Generalization through High-Similarity Data

### v2

# Bingxiang He1 , Ning Ding [^1] footnotemark: 1 , Cheng Qian1,211footnotemark: 1  , Jia Deng3, Ganqu Cui1, Lifan Yuan2 Haiwen Hong4, Huan-ang Gao1, Longtao Huang4, Hui Xue4, Huimin Chen1 Zhiyuan Liu1 , Maosong Sun122footnotemark: 2

@@ Abstract
Understanding alignment techniques begins with comprehending zero-shot generalization brought by instruction tuning, but little of the mechanism has been understood. Existing work has largely been confined to the task level, without considering that tasks are artificially defined and, to LLMs, merely consist of tokens and representations. To bridge this gap, we investigate zero-shot generalization from the perspective of the data itself. We first demonstrate that zero-shot generalization happens very early during instruction tuning, with loss serving as a stable indicator. Next, we investigate training data arrangement through similarity and granularity perspectives, confirming that the timing of exposure to certain training examples may greatly facilitate generalization on unseen tasks. Finally, we propose a more grounded training data arrangement framework, Test-centric Multi-turn Arrangement, and show its effectiveness in promoting continual learning and further loss reduction. For the first time, we show that zero-shot generalization during instruction tuning is a form of similarity-based generalization between training and test data at the instance level. [^1] Our code is released at https://github.com/thunlp/Dynamics-of-Zero-Shot-Generalization.
[…]
footnote
[Uncaptioned image] The Right Time Matters: Data Arrangement Affects Zero-Shot Generalization in Instruction Tuning
Bingxiang He1thanks: Equal contribtuion. , Ning Ding [^1] footnotemark: 1 , Cheng Qian1,211footnotemark: 1  , Jia Deng3, Ganqu Cui1, Lifan Yuan2 Haiwen Hong4, Huan-ang Gao1, Longtao Huang4, Hui Xue4, Huimin Chen1 Zhiyuan Liu1thanks: Corresponding Author. , Maosong Sun122footnotemark: 2 1Tsinghua University,  2University of Illinois Urbana-Champaign, 3Renmin University of China,  4Alibaba Group hebx24@mails.tsinghua.edu.cndn97@mail.tsinghua.edu.cn chengq9@illinois.edu
@@ 1 Introduction
The extraordinariness of large language models (LLMs) was originally brought by the zero-shot generalization of instruction tuning (Brown et al., 2020). Early studies have found that when diverse prompts are added to the inputs of traditional natural language processing (NLP) tasks and fed into the model for instruction tuning, the model can generalize to tasks it has never encountered before (Chung et al., 2024; Longpre et al., 2023; Sanh et al., 2021; Wang et al., 2022; Wei et al., 2021). To date, instruction tuning (Chung et al., 2024; Sanh et al., 2021; Wei et al., 2021) has become a crucial phase in LLM training, often preceding methods that incorporate preference data. In the meantime, the concept of “task” is also becoming increasingly blurred. Researchers are no longer constructing instruction data in the ways traditional NLP tasks dictate, but rather, they hope these tasks will be as close to reality and as diverse as possible (Chiang et al., 2023; Ding et al., 2023; Rajani et al., 2023; Taori et al., 2023; Zhao et al., 2024b).
@@
Although nearly all LLMs benefit from the zero-shot generalization brought about by instruction tuning, the in-depth and fine-grained research on this phenomenon is still insufficient. Particularly, there are few accurate and comprehensive conclusions about when and in what form it occurs, and how it would be influenced at an instance level. Existing approaches (Song et al., 2019; Vu et al., 2020; Zamir et al., 2018; Kim et al., 2023; Muennighoff et al., 2023; Zhou et al., 2022; Lee et al., 2024) assume that human-defined “tasks” and even “categories” are sufficiently reasonable, but this is often not the case, as gaps exist regarding how humans and LLMs perceive the instruction tuning data. To this end, we strive to break free from the task-level framework and explore “generalization” from a more granular temporal perspective.
@@
In this paper, we conduct a comprehensive investigation of the zero-shot generalization during instruction tuning and attempt to answer some critical questions: In instruction tuning, i) when does zero-shot generalization occur? ii) how can we more accurately understand the role of data in zero-shot generalization? iii) how can we effectively improve zero-shot generalization?
To answer the first research question, we attempt to pinpoint the timing of zero-shot generalization during instruction tuning. We find that significant generalization occurs after training on just 160 examples, demonstrating that instruction-following capabilities can be unlocked with minimal data, similar to existing works (Chen et al., 2023; Gudibande et al., 2023; Zhao et al., 2024a; Zhou et al., 2024) but we provide a more granular analysis. Our analysis shows that loss serves as a more reliable indicator of zero-shot generalization compared to traditional evaluation metrics like ROUGE and Exact-Match. (Section 3)
For the second question, we track LLaMA-2-7B’s loss on 255 unseen test tasks while instruction-tuning on 1,600 training tasks. We find that simply reordering training data produces diverse test loss patterns—from sudden improvements to gradual changes or fluctuations. This indicates that while the model can achieve rapid generalization under certain conditions, it may also show delayed generalization or fail to generalize depending on the training data arrangement. More importantly, these observations indicate that the timing of exposure to certain training examples may greatly facilitate generalization on unseen tasks, highlighting the crucial role of data arrangement in effective instruction tuning.
@@
To understand the role of data in this process, we identify two perspectives: similarity and granularity. From a similarity perspective, we discover that the model’s generalization is not truly “zero-shot”, as a high resemblance between the training and test data distributions could significantly impact generalization. Using combined cosine similarity measures, we find that test loss drops suddenly when the model encounters training examples highly similar to the test set, particularly when such examples appear early in training. From a granularity perspective, the artificially defined “tasks” are not suitable for measuring generalization. Instead, the similarity measure at the instance level serves as a better and more essential indicator. We reveal through experiments that, by treating all data points equally without the constraints of pre-defined tasks, we can better improve zero-shot generalization. As shown in LABEL:fig:intro, encountering highly similar and fine-grained training data earlier during instruction tuning enables better generalization. (Section 4)
While the combined similarity measures we adopt reveal the role of training data, they have inherent limitations, as they treat the test set holistically, ignoring the role of individual test data points. To address this, we propose the Test-centric Multi-turn Arrangement (TMA) framework, which organizes training data in a test-centric way, even without predefined tasks. Through the lens of TMA, we reveal that accessing highly similar data during instruction tuning promotes continual learning and further loss reduction. (Section 5)
@@
  • We show that zero-shot generalization occurs at the very early stage during instruction tuning, while loss serves as a more stable and fair metric compared to traditional ones.
  • We emphasize the critical role of timing in data exposure as a key perspective to gain a deeper understanding of zero-shot generalization, revealing that encountering highly similar and fine-grained training data earlier during instruction tuning enables better generalization.
  • We propose the Test-centric Multi-turn Arrangement framework, and show that accessing high-similarity data during instruction tuning can facilitate continual learning and further loss reduction.
2 Related Work
LLMs have been proven capable of zero-shot generalization across a variety of downstream tasks (Brown et al., 2020), and instruction tuning has emerged as the most effective method to achieve this (Chung et al., 2024; Sanh et al., 2021; Wei et al., 2021). The zero-shot generalization phenomenon resulting from instruction tuning is crucial for building general LLMs. Multi-task datasets designed for this purpose have continuously iterated in quality, quantity, and diversity, and numerous studies have explored how zero-shot generalization occurs during the instruction tuning process. Sanh et al. (2021) constructed P3 using explicit multitask learning, demonstrating that explicit task prompt templates can promote zero-shot generalization. Wang et al. (2022) created the Super Natural Instructions V2 (NIV2) dataset, which comprises over 1600 task types, and empirically showed that more observed tasks, an adequate number of training instances, and larger models improve generalization. Meta introduced OPT-IML (Iyer et al., 2022), investigating the impacts of dataset scale and diversity, different task sampling strategies, and the presence of demonstrations on generalization. Subsequently, Longpre et al. (2023) proposed the Flan Collection, which encompasses up to 1836 tasks, and pointed out that scaling the number of tasks and model size, as well as incorporating chain-of-thought data, can dramatically improve performance on unseen tasks.
In addition, a line of research focuses on understanding the relationships in task-pair transfer (Song et al., 2019; Vu et al., 2020; Zamir et al., 2018), suggesting that not all tasks contribute positively to zero-shot generalization; some tasks may even result in negative transfer effects (Kim et al., 2023; Muennighoff et al., 2023; Zhou et al., 2022). However, a significant limitation of the aforementioned works is that, whether training on one task and then evaluating on another (Kim et al., 2023; Zhou et al., 2022), or simply calculating intermediate task (Vu et al., 2020) or instruction (Lee et al., 2024) transfer scores, all these efforts to select the most informative tasks to promote zero-shot generalization are confined within the “task” framework. This approach is based on a premise: the human-defined “tasks” and even “categories” are sufficiently reasonable. This is precisely the issue our study strives to address: breaking free from the task-level framework to explore “generalization” at a more fundamental level.
@@ 3 Positioning Zero-Shot Generalization
Early research shows that instruction tuning, which applies to various NLP tasks formatted with instructions, can generalize to various unseen tasks. However, most studies (Chung et al., 2024; Iyer et al., 2022; Longpre et al., 2023) focus on integrating diverse tasks or instruction templates, using human-generated or synthetic data, or exploring different fine-tuning strategies, while few studies address the timing of zero-shot generalization. To bridge this gap, we first seek to identify when zero-shot generalization actually occurs during instruction tuning, and then justify that loss serves as a more stable and fair metric to measure zero-shot generalization compared to traditional ones including ROUGE series, Exact-Match, Reward Model scores, etc. We begin by giving a formalization of zero-shot generalization.
Formalization. In multi-task scenarios, zero-shot generalization refers to the ability to perform effectively on unseen tasks ($\mathcal{T}_{\text{Unseen}}$), while only trained on a subset of tasks ($\mathcal{T}_{\text{Seen}}$). For each task $T\in\mathcal{T}_{\text{Seen}}\cup\mathcal{T}_{\text{Unseen}}$, there exists an instructional description $I_{T}$, as well as several instances, where each instance is composed of an input $x_{T,i}$ and an output $y_{T,i}$. We define a model $M$ as capable of generalization on unseen tasks if, after training on every task in $\mathcal{T}_{\text{Seen}}$, given an unseen task $T\in\mathcal{T}_{\text{Unseen}}$, and for any $(x_{T,i},y_{T,i})$, the model’s output $\hat{y}=M(I_{T},x_{T,i})$ and the label $y_{T,i}$ achieve a score surpassing a certain threshold, with regard to the selected metrics, indicating successful generalization on the task $T$.
3.1 Early Zero-Shot Generalization
@@
Data and Settings. We utilize three multi-task datasets, namely Natural Instructions V2 (NIV2) (Wang et al., 2022), Public Pool of Prompts (P3) (Sanh et al., 2021), and Flan-mini (Ghosal et al., 2023), for our analysis. For NIV2, we utilize the default track and training-test split for instruction tuning and evaluation. For P3, we employ training and test tasks consistent with the vanilla T0 model [^2] https://huggingface.co/bigscience/T0pp. For Flan-mini, we randomly partition the training and test tasks. We choose pre-trained LLaMA-2-7B (Touvron et al., 2023) as our base model. For other details including dataset, hyper-parameters, and prompt template, please refer to Appendix A. All subsequent experiments are based on this setup, where we adopt to save a series of full-parameter fine-tuning checkpoints and evaluate each on the unseen test set to observe the results regarding specified metrics.
Metrics. We experiment with multiple metrics, including Exact-Match, ROUGE-1, ROUGE-L, and RM score, to test their ability to reasonably reflect zero-shot generalization. For P3 and Flan-mini, Exact-Match is commonly applied in previous studies due to its simplicity, while NIV2 additionally incorporates ROUGE series as metrics. Besides, in reinforcement learning scenarios, the reward model (RM) often plays a vital role (Cui et al., 2023; Yuan et al., 2024) and serves as a proxy for human preferences. This makes the RM score also a plausible metric to reflect zero-shot generalization. Empirically, we use UltraRM-13B (Cui et al., 2023) as the reward model when measuring RM score.
Results. We demonstrate that zero-shot generalization occurs at a very early stage during instruction tuning regardless of metric choice. As depicted in the left plot of Figure 1, using ROUGE series and Exact-Match as metrics, the scores rise from approximately 15 to over 35 in merely 10 training steps, indicating significant generalization with only 160 training samples in our setting. In the middle plot, the RM score exhibits a similar trend, stabilizing around 50 steps across all three datasets.
Despite the similar trend they present, it should be noted that ROUGE-1, ROUGE-L, and Exact-Match as metrics all entail the resulting curves being seriously unstable, while the RM score for NIV2 is significantly higher than those for the other two datasets, indicating a certain bias induced (more details discussed in  Section A.5). This leads us to seek a more reasonable metric as an indicator to evaluate zero-shot generalization.
3.2 Loss as Generalization Indicator
@@
Data and Settings. We use the same dataset as in the previous experiment and generate outputs for sampled test data points using a series of instruction-tuned checkpoints we derived. We then calculate the average cross-entropy loss against the corresponding labels within each step. Please refer to Section A.4 for more details.
Results. Zero-shot generalization similarly occurs at an early stage of instruction tuning with loss as the metric. As shown in the right plot of Figure 1, all three datasets reach their lowest points in terms of loss within less than 50 steps, which strengthens the conclusion that generalization occurs early.
@@
4 Data Arrangement Effects on Zero-Shot Generalization
Acknowledging the importance of metrics in measuring the positioning of zero-shot generalization, we next seek to investigate why generalization occurs at an early stage and what role training data plays during this phase. Our initial focus lies in the analysis of how various simple training data arrangements affect zero-shot generalization in a fine-grained manner. Then, we investigate the facilitation of zero-shot generalization from both data similarity and granularity perspectives.
4.1 Pilot Study
@@
The model receives only a limited amount of data at the early stage of instruction tuning. Therefore, despite the scarcity, these data ought to play a significant role in facilitating generalization. Guided by this intuition, we conduct a pilot study to explore the impact of exposure to different training data arrangements from a temporal perspective.
Data and Settings. We apply 1600 Flan-mini training tasks to get a series of instruction-tuned checkpoints and evaluate them on various unseen test tasks in Flan-mini. As shown in LABEL:fig:exp_settings_1, we examine the following three training data arrangements:
  • Round Robin: We select one data point from each task to form a data batch, ensuring that training tasks are evenly distributed.
  • Cluster: We arrange all data from each task together, resulting in task-level clusters throughout the entire training dataset.
@@
Results. Training data arrangements lead to distinct loss curve patterns. As shown in Figure 2, random and round-robin scheduling produce similar patterns, with round-robin being an extreme form of shuffling. In contrast, cluster scheduling shows significant differences, including sudden drops in average loss at specific steps during instruction tuning. This demonstrates that leveraging a small subset of data can induce substantial loss reduction, and the same test task may exhibit varying generalization behaviors under different data arrangements.
4.2 Through Data Similarity and Granularity
We have observed that training data arrangements could lead to significant changes in the loss curve and that the timing of presence for certain data may greatly facilitate generalization on unseen tasks. With these findings, we naturally ask what is the best arrangement of training data, and how to arrange these “certain data” that improve early generalization. In the following subsections, we seek to address these questions through two perspectives: similarity and granularity.
4.2.1 Effect of High-Similarity Data